OpenStudy (anonymous):

can someone help me with how to find the equation of the perpendicular line in slope intercept form containing the point (-3,7) where the equation is 5x+3y=8

OpenStudy (anonymous):

first you need the slope of \(5x+3y=8\) do you know what it is?

OpenStudy (anonymous):

I think I figured it out - I got -5/3

OpenStudy (anonymous):

so I know had to flip it - this is what I was working with - but don't think I am doing it right, y-7=3/5(x-5)